FAA
23: the measure of the complement of the angle with measure 45 degrees.

Answer by checkley75(3666)

COMPLEMENTARY ANGLES ADD UP TO 90 DEGREES.
THUS 45+X=90 0R X=90-45 OR X=45 DEGREES FOR THE COMPLEMENTARY ANGLE TO A 45 DEGREE ANGLE.